We are given the lengths of the three altitudes in a triangle.And the length of any one side is asked


Wht is the procedure???

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Palas

let altitude are x,y,z

we know  area of triangle Δ =ax/2 = by/2 = cz/2

calculate a,b,c in terms of x,y,z and Δ

now calculate s,s-a,s-b,s-c  in terms of x,y,z,Δ

 

now use the hero's formula of triangle to calculate area.

and put the value of area in above equation to find a,b,c

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin